A jet can travel the 6000 km distance between Washington, D.C. and London in 6 hours with the wind. The return trip against the same wind takes 7.5 hours. Find the rate of the jet in still air and the rate of the wind.
Answer by oberobic(2304)

d = r*t is the basic distance equation
d = 6000 km
.
t with the tail wind = 6 hr
r with the tail wind = speed of the plane + wind speed = s + w
.
t with the head wind = 7.5 hr
r with the head wind = speed of the plane - wind speed = s-w
.
(s+w)*6 = 6000
(s-w)*7.5 = 6000
.
s + w = 1000
s - w = 800
-------------
2s = 1800
s = 900 km/h
.
s + w = 1000
w = 100
.
Check the anwer by calculating the return trip.
.
(900-100) * 7.5 = 800 * 7.5
800 * 7.5 = 6000 km
.
Answer: The rate of the jet in still air is 900 km/h. The rate of the wind is 100 km/hr.
.
Done.